a right-angled triangle has an area of 20cm^2.
The width of the triangle is n+5cm.
The height is 6cm less than the width.
FInd the value of n

Answer by Cromlix(4381)

A right angled triangle.
Width = (n + 5)
Height = (n + 5) - 6
Height = (n - 1)
Area of a triangle = 1/2 (base x height)
20 = 1/2(n - 1)x(n + 5)
40 = n^2 + 4n - 5
n^2 + 4n - 45 = 0
(n - 5)(n + 9) = 0
n + 9 = 0
n = -9 (no answer, as negative value)
n - 5 = 0
n = 5
This would make the width = 10 cm
and the height = 4 cm
Area = 1/2(10 x 4) = 20 cm^2
Hope this helps.
:-)